OpenStudy (anonymous):

. Solve the equation. Check for extraneous solutions. |4x+3|=9+2x

OpenStudy (anonymous):

@triciaal @KlOwNlOvE

OpenStudy (anonymous):

4x+3=9+2x 4x+3=-9-2x get that 2x over with 4

OpenStudy (triciaal):

2 situations -(4x + 3) = 9 + 2x and 4x + 3 = 9 + 2x

OpenStudy (triciaal):

-4x -3 = 9 + 2x -6x = 12 x = -2 2x = 6 x = 3

OpenStudy (triciaal):

check in original when x = 3 4(3) + 3 = 12 + 3 = 15 9 + 2(3) =9 + 6 = 15 when x = -2 4(-2) + 3 = -8 + 3 = -5 absolute value = 5 9 + 2(-2) = 9 -4 = 5

OpenStudy (triciaal):

any questions?

OpenStudy (anonymous):

yes, how would I fill this in x = _____ or _____

OpenStudy (anonymous):

@triciaal

OpenStudy (triciaal):

what are the 2 values for x?

OpenStudy (triciaal):

read the posts there were 2 situations we have 2 values for x that satisfy the equation